A Complete Bibliography of Digital Threats: Research and Practice (DTRAP)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

02 March 2022 Version 1.03

Title word cross-reference

.eu [DSV^{+}21].

1 [PLD^{+}20]. 11510 [Koo20].

ABI [AVS^{+}22]. Access [OMHY20]. Accurate [SMK20]. ACSAC’19 [PLD^{+}20, PLD^{+}21]. ACSAC’20 [PLS22]. Activities [LLM20]. Advanced [IKT^{+}21, LLM20]. Advertising [CYGL20]. Aegis [SBU21]. Affordable [DD20]. against [DKP21, OMHY20]. AI [DWN21]. Analysis [KCH^{+}20, RUPT22]. Analysts [HAH^{+}21]. Analyze [TCBK20]. Android [GTL^{+}20, KCH^{+}20]. Annotation [KPBZ21]. Anonymization [AYP^{+}21].

Antivirus [GLS21]. API [UOI^{+}21]. Appliances [WMY20]. Applications [OMHY20]. Apps [GTL^{+}20]. ARIMA [KB20]. Artifacts [DIPQ22]. Artificial [WASC22]. Assessing [HAH^{+}21, LRMZA21]. Assets [SLBC21]. Assisted [KIP21]. Assurance [KB20]. ATFuzzer [KCH^{+}20]. Attack [DD20, SKM^{+}20, Sun22]. Attacks [AVS^{+}22, Bur20, DKP21, OMHY20, PMP20]. Audit [GSS21]. Authentication [KIP21, Sun22]. Automated [KPBZ21, SMK20]. Automatic [UOI^{+}21]. Automation [BSM^{+}22]. aware [MBP22, SBU21]. Awareness [APY21, DRS^{+}21, GDP20, KB20, SLBC21]. AZSecure [SLBC21].

Based [DD20, AYP^{+}21, Bur20, DKP21, KIP21, RVD^{+}21]. Behavioural [IKT^{+}21]. Benchmark [KPBZ21]. Better [Sac20, Van21]. Beyond [VS20]. Biases [DWN21]. Binaries [UOI^{+}21]. Binary [ADJ^{+}20, DIPQ22, FRF22]. Biometrics [IKT^{+}21, SKM^{+}20]. Blockchain [Ami21, MMT^{+}21]. Blocklisting [MS21]. Bots [IKT^{+}21]. Building [UOI^{+}21].

Card [GSS21]. Card-style [GSS21]. Centers [RCGA22]. Centrality [APY21]. Certain [Bur20]. Certificate [UKM20]. Challenges [RUPT22]. Channel [DD20, OMHY20]. Character [MBP22]. Characterization [Bur20]. Chat [OASD^{+}22]. Chosen [Sun22]. Claim [KPBZ21]. Code [SRBS20]. Collaboratively [LRMZA21]. Combining [IKT^{+}21]. Commodity [CPDH20]. Condensation [AYP^{+}21]. Condensation-based [AYP^{+}21]. Connect [Koo20]. Consistent [KPBZ21]. Conspicuous [DIPQ22]. Context [SBU21, ST20, VSV^{+}21]. Context-aware [SBU21]. Context-sensitive [ST20]. Continuous [Sac20]. Control [DKP21, GLS21]. Control-based [DKP21]. Cookies [SRBS20]. Coping [VSV^{+}21]. Correlation [SKM^{+}20]. Countermeasures [CYGL20, FRF22]. Credible [HZB21]. Cues [VSV^{+}21]. Cut [GLS21]. Cut-and-Mouse [GLS21]. CVE [Koo20]. CVE-2019-11510 [Koo20]. Cyber [APY21, BSM^{+}22, SLBC21]. Cybersecurity [GDP20, LRMZA21, Van21].

Dark [SLBC21]. Data [AYP^{+}21, BSM^{+}22, GSS21, MBP22, PMP20, SDA^{+}22, dWBvdH22]. Data-only [PMP20]. Database [ST20]. DDoS [Bur20]. Debloating [ADJ^{+}20]. Deception [HBRV^{+}21]. Decision [HAH^{+}21]. Decrease [GSS21]. Deep [RVD^{+}21]. Defences [HBRV^{+}21]. Delivery [LRMZA21]. Detect [FRF22]. Detection [DKP21, IKT^{+}21, KPBZ21, MS21, RVD^{+}21, SMK20, ST20, VS20, WASC22, ZYW22, ZJPZ20, dWBvdH22]. Developing [KPBZ21]. Differential [AYP^{+}21]. Digital [LM20, SMK20]. Directional [APY21]. Dirichlet [ZYW22]. Dissected [ZBF22]. DNS [Bur20, NPT21]. DNS-based [Bur20]. DNS-over-TLS [NPT21]. Documents [SMK20]. Domain [DSV^{+}21, HZB21, MS21]. driven [ZJPZ20]. Dynamic [DIPQ22, FRF22, KCH^{+}20, dWBvdH22].

E-Voting [MMT^{+}21]. Early

[LLM20, ZJPZ20]. Ecosystem [MS21]. Editorial [LM20, Met21]. EEG [SKM^{+}20]. Election [Met21]. ELF [TCBK20]. Emerging [KIP21]. Enabling [BSM^{+}22]. Encrypted [MBP22]. Energy [LRMZA21]. Engine [UOI^{+}21]. Engineering [UOI^{+}21]. Enterprise [SDA^{+}22, WMY20]. Environments [AVS^{+}22]. EPSS [JRE^{+}21]. Evaluating [DIPQ22]. Evaluations [DD20]. Evasion [FRF22]. Exchange [BSM^{+}22]. Execution [AVS^{+}22]. Experimenting [RCGA22]. Experts [HZB21]. Exploit [JRE^{+}21]. Exploitation [GTL^{+}20]. Exploiting [GLS21]. ExSol [LRMZA21]. Extended [UKM20].

Factchecking [KPBZ21]. Fake

[VSV^{+}21, ZLPA21, ZJPZ20]. False [Sac20]. Faulty [AVS^{+}22]. Features [DIPQ22, TCBK20]. Field [Koo20]. Financial [HZB21]. Fingerpointing [Sac20]. FIRST [CV20]. Foundations [ZBF22]. Framework [KCH^{+}20, SBU21, TCBK20]. Frameworks [FRF22].

Gait [KIP21]. Gaps [GDP20]. Generation [MS21]. Ghost [GLS21]. Graybox [ST20]. Group [OASD^{+}22].

Hacker [SLBC21]. Hawkes [ZYW22]. History [SRBS20]. Home [SBU21]. Horizon [VS20]. Host [RVD^{+}21]. Host-based [RVD^{+}21].

Identification [LLM20]. Identifying [HZB21]. Improvement [Sac20]. Inaugural [LM20]. independent [SBU21]. Inference [SKM^{+}20, SDA^{+}22]. Informing [SLBC21]. Input [SDA^{+}22]. Inputs [GLS21]. Inroduction [ZLPA21]. Insider [Ano22, WASC22, ZYW22]. Insiders [ZBF22]. Instrumentation [DIPQ22, FRF22]. Integrate [Sac20]. Intelligence [BSM^{+}22, SLBC21]. Interception [ST20, WMY20]. Interface [KCH^{+}20]. Introduction [Ano22, CV20, DRS^{+}21, PLD^{+}20, PLD^{+}21, PLS22, RUPT22]. Intrusion [CPDH20, DKP21, RVD^{+}21]. Investigating [BSM^{+}22]. Investigation [VSV^{+}21]. Invoking [WASC22]. IoT [TCBK20]. Issue [Ano22, CV20, LM20, Met21, PLD^{+}20, PLD^{+}21, PLS22, RUPT22].

Knowledge [BSM^{+}22].

Laplacian [APY21]. Large [ADJ^{+}20]. Large-scale [ADJ^{+}20]. Leakage [DD20]. Leakage-Based [DD20]. Learning [Bur20, RVD^{+}21, SMK20, TCBK20, dWBvdH22]. Less [GSS21]. Libraries [ADJ^{+}20]. limiting [GSS21]. Logs [IKT^{+}21]. Looking [VS20].

Machine [TCBK20, dWBvdH22]. Malicious [DSV^{+}21, SMK20]. Malware [Bur20, LLM20, RUPT22, TCBK20, dWBvdH22]. Marked [ZYW22]. Matching [MBP22]. Measure [WASC22]. Media [DWN21]. Memory [PMP20]. Memory-resident [PMP20]. Methodology [RCGA22]. Methods [RVD^{+}21]. Metrics [KB20]. Mitigating [PMP20]. Mitigation [WASC22]. Mobile [dWBvdH22]. Model [BSM^{+}22, SMK20, SKM^{+}20, ZJPZ20]. Modeling [SDA^{+}22]. Monitoring [Sac20]. Mouse [GLS21, IKT^{+}21].

Name [DSV^{+}21]. Nature [DRS^{+}21]. Netherlands [Koo20]. Network [AYP^{+}21, HBRV^{+}21]. Networks [WASC22]. Neural [WASC22]. News [DWN21, VSV^{+}21, ZLPA21, ZJPZ20]. Note [Koo20].

Online [CYGL20]. only [PMP20]. Opcode [TCBK20]. Operating [CPDH20]. Operation [RCGA22]. Operational [DSV^{+}21]. Opportunities [GDP20]. Organizational [SDA^{+}22]. Outsourced [MBP22].

Part [PLD^{+}20]. Past [SRBS20]. Pattern [MBP22]. PDoT [NPT21]. People [UKM20]. Perceptions [OASD^{+}22, UKM20]. Perfect [DWN21]. Persistent [LLM20]. Perspectives [ZLPA21]. Pitaya [DD20]. Platform [SBU21]. Platform-independent [SBU21]. Point [AVS^{+}22]. Poisoning [AVS^{+}22]. Portal [SLBC21]. Positives [Sac20]. Power [DD20]. Practical [DD20]. Practice [Ami21, LM20]. Prediction [JRE^{+}21]. Premadoma [DSV^{+}21]. Prevent [DSV^{+}21]. Privacy [AYP^{+}21, MBP22, OASD^{+}22]. Privacy-aware [MBP22]. Private [NPT21]. Proactive [VS20]. Processes [ZYW22]. Processing [SRBS20]. Proposing [BSM^{+}22]. Protected [DKP21]. Protecting [LRMZA21, PMP20]. Protocol [Sun22]. Psychological [DWN21]. Pulse [Koo20].

Quality [KB20].

Random [Sun22]. Real [HZB21]. Real-world [HZB21]. Redacted [SDA^{+}22]. Registrations [DSV^{+}21]. Representation [BSM^{+}22]. Request [SRBS20]. Research [LM20, ZLPA21, ZBF22]. resident [PMP20]. Resource [OMHY20]. Results [WASC22]. Reverse [UOI^{+}21]. Risk [GSS21]. Risk-limiting [GSS21]. Risks [LRMZA21]. Robotic [DKP21].

SAE [Sun22]. SAIBERSOC [RCGA22]. Sample [GSS21]. SCA [DD20]. SCA-Pitaya [DD20]. scale [ADJ^{+}20]. Schema [KPBZ21]. Scoring [JRE^{+}21]. Script [UOI^{+}21]. Secure [Koo20, MMT^{+}21]. Securing [OMHY20]. Security [KB20, Met21, OASD^{+}22, RCGA22, Sac20, SBU21, WMY20]. Sensitive [PMP20, ST20]. Sensor [KIP21]. Sensor-based [KIP21]. Server [SRBS20]. Server-side [SRBS20]. Setup [DD20]. Shared [ADJ^{+}20]. Sharing [BSM^{+}22, VSV^{+}21]. Side [DD20, OMHY20, SRBS20]. Side-Channel [DD20, OMHY20]. Simulating [SDA^{+}22]. Situational [APY21, DRS^{+}21, GDP20, KB20, SLBC21]. Sizes [GSS21]. Smart [SBU21]. Smartphone [dWBvdH22]. Smartphones [KCH^{+}20]. Sniffing [SRBS20]. SOC [HAH^{+}21]. Social [DWN21]. Software [GLS21]. Solution [DSV^{+}21]. Sorry [WMY20]. Special [Ano22, CV20, Met21, PLD^{+}20, PLD^{+}21, PLS22, RUPT22]. Spoofing [KIP21]. SSL [Koo20]. SSL-VPN [Koo20]. State [WMY20]. Stealthy [DKP21]. Stored [ST20]. Storm [DWN21]. Style [GSS21]. Supplemented [KB20]. Support [HAH^{+}21, NPT21]. Survey [CYGL20]. Survivability [CPDH20]. Synthesized [GLS21]. System [JRE^{+}21]. Systematisation [ZBF22]. Systems [CPDH20, DIPQ22, LRMZA21, SBU21].

TAGS [KIP21]. Technical [CYGL20]. Techniques [Bur20, DKP21, FRF22]. TEE [NPT21]. Theory [ZJPZ20]. Theory-driven [ZJPZ20]. Thinking [SKM^{+}20]. Thoughts [VS20]. Threat [BSM^{+}22, KIP21, LLM20, SLBC21, VSV^{+}21, WASC22, ZYW22]. Threats [Ano22, CYGL20, LM20, VS20]. Thwart [GTL^{+}20]. Timing [SRBS20]. TLD [DSV^{+}21]. TLS [NPT21, UKM20, WMY20]. Tool [HAH^{+}21, RCGA22]. Tracers [UOI^{+}21]. Traces [AYP^{+}21]. Traffic [ST20]. Treadmill [KIP21]. Trends [RUPT22, ZLPA21]. Triage [LLM20]. Trust [UKM20]. Trusted [AVS^{+}22].

Understanding [Bur20, Van21]. Unit [AVS^{+}22]. Unpredictability [HBRV^{+}21]. Unsupervised [Bur20]. Unveiled [SKM^{+}20]. User [OASD^{+}22]. Using [HBRV^{+}21, TCBK20, ZYW22, dWBvdH22].

Value [Sun22]. Vehicles [DKP21]. Version [UKM20]. Veto [OMHY20]. Vetting [GTL^{+}20]. Virus [VSV^{+}21]. Vol [PLD^{+}21]. Voting [MMT^{+}21]. VPN [Koo20]. Vulnerabilities [Ami21, GTL^{+}20]. Vulvet [GTL^{+}20].

Wearable [KIP21]. Web

[IKT^{+}21, SLBC21]. Will [UKM20]. Work [GSS21]. Working [UKM20]. world [HZB21]. WPA3 [Sun22].

XSS [ST20].

Zika [VSV^{+}21].

References

Agadakos:2020:LSD

[ADJ^{+}20] Ioannis Agadakos, Nicholas Demarinis, Di Jin, Kent Williams-King, Jearson Alfajardo, Benjamin Shteinfeld, David Williams-King, Vasileios P. Kemerlis, and Georgios Portokalidis. Largescale debloating of binary shared libraries. Digital Threats: Research and Practice (DTRAP), 1(4):19:1–19:28, December 2020. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3414997.

Amiet:2021:BVP

[Ami21] Nils Amiet. Blockchain vulnerabilities in practice. Digital Threats: Research and Practice (DTRAP), 2(2):8:1–8:7, March 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3407230.

Anonymous:2022:ISI

[Ano22] Anonymous. Introduction to the special issue on insider threats. Digital Threats: Research and Practice (DTRAP), 3(1):1:1–1:3, March 2022. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3477501.

Aksoy:2021:DLC

[APY21] Sinan G. Aksoy, Emilie Purvine, and Stephen J. Young. Directional Laplacian centrality for cyber situational awareness. Digital Threats: Research and Practice (DTRAP), 2(4):28:1– 28:28, December 2021. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3450286.

Alder:2022:FPU

[AVS^{+}22] Fritz Alder, Jo Van Bulck, Jesse Spielman, David Oswald, and Frank Piessens. Faulty point unit: ABI poisoning attacks on trusted execution environments. Digital Threats: Research and Practice (DTRAP), 3(2):13:1–13:26, June 2022. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3491264.

Aleroud:2021:ANT

[AYP^{+}21] Ahmed Aleroud, Fan Yang, Sai Chaithanya Pallaprolu, Zhiyuan Chen, and George Karabatis. Anonymization of network traces data through condensation-based differential privacy. Digital Threats: Research and Practice (DTRAP), 2(4):30:1–30:23, December 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3425401.

Bromander:2022:ISC

[BSM^{+}22] Siri Bromander, Morton Swimmer, Lilly Pijnenburg Muller, Audun Jøsang, Martin Eian, Geir Skjøtskift, and Fredrik Borg. Investigating sharing of cyber threat intelligence and proposing a new data model for enabling automation in knowledge representation and exchange. Digital Threats: Research and Practice (DTRAP), 3(1):6:1–6:22, March 2022. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3458027.

Burton:2020:ULT

[Bur20] Renée Burton. Unsupervised learning techniques for malware characterization: Understanding certain DNS-based DDoS attacks. Digital Threats: Research and Practice (DTRAP), 1(3):14:1–14:26, September 2020. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3377869.

Chevalier:2020:ISC

[CPDH20] Ronny Chevalier, David Plaquin, Chris Dalton, and Guillaume Hiet. Intrusion survivability for commodity operating systems. Digital Threats: Research and Practice (DTRAP), 1(4):21:1– 21:30, December 2020. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3419471.

Cormack:2020:IFS

[CV20] Andrew Cormack and Jeroen Van der Ham. Introduction to the FIRST special issue. Digital Threats: Research and Practice (DTRAP), 1(1): 2:1–2:3, March 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3380746.

Cai:2020:TOA

[CYGL20] Yegui Cai, George O. M. Yee, Yuan Xiang Gu, and Chung-Horng Lung. Threats to online advertising and countermeasures: a technical survey. Digital Threats: Research and Practice (DTRAP), 1(2): 11:1–11:27, July 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3374136.

Durvaux:2020:SPP

[DD20] François Durvaux and Marc Durvaux. SCA-Pitaya: a practical and affordable sidechannel attack setup for power leakage-based evaluations. Digital Threats: Research and Practice (DTRAP), 1(1):3:1– 3:16, March 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3371393.

DElia:2022:EDB

[DIPQ22] Daniele Cono D’Elia, Lorenzo Invidia, Federico Palmaro, and Leonardo Querzoni. Evaluating dynamic binary instrumentation systems for conspicuous features and artifacts. Digital Threats: Research and Practice (DTRAP), 3(2):10:1– 10:13, June 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3478520.

Dash:2021:SAA

[DKP21] Pritam Dash, Mehdi Karimibiuki, and Karthik Pattabiraman. Stealthy attacks against robotic vehicles protected by control-based intrusion detection techniques. Digital Threats: Research and Practice (DTRAP), 2(1):7:1–7:25, March 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3419474.

Dykstra:2021:INS

[DRS^{+}21] Josiah Dykstra, Neil Rowe, Timothy Shimeall, Angela Horneman, and Marisa Midler. Introduction: On the nature of situational awareness. Digital Threats: Research and Practice (DTRAP), 2(4):25:1– 25:3, December 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3462334.

Desmet:2021:POS

[DSV^{+}21] Lieven Desmet, Jan Spooren, Thomas Vissers, Peter Janssen, and Wouter Joosen. Premadoma: an operational solution to prevent malicious domain name registrations in the .eu TLD. Digital Threats: Research and Practice (DTRAP), 2(1):2:1–2:24, March 2021. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3419476.

deWit:2022:DDM

[dWBvdH22] J. S. Panman de Wit, D. Bucur, and J. van der Ham. Dynamic detection of mobile malware using smartphone data and machine learning. Digital Threats: Research and Practice (DTRAP), 3(2):9:1–9:24, June 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3484246.

Datta:2021:PSS

[DWN21] Pratim Datta, Mark Whitmore, and Joseph K. Nwankpa. A perfect storm: Social media news, psychological biases, and AI. Digital Threats: Research and Practice (DTRAP), 2(2): 15:1–15:21, June 2021. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3428157.

Filho:2022:ECT

[FRF22] Ailton Santos Filho, Ricardo J. Rodríguez, and Eduardo L. Feitosa. Evasion and countermeasures techniques to detect dynamic binary instrumentation frameworks. Digital Threats: Research and Practice (DTRAP), 3(2):11:1– 11:28, June 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3480463.

Gutzwiller:2020:GOS

[GDP20] Robert Gutzwiller, Josiah Dykstra, and Bryan Payne. Gaps and opportunities in situational awareness for cybersecurity. Digital Threats: Research and Practice (DTRAP), 1(3):18:1– 18:6, September 2020. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3384471.

Genc:2021:CMG

[GLS21] Ziya Alper Genç, Gabriele Lenzini, and Daniele Sgandurra. Cut-and-mouse and ghost control: Exploiting antivirus software with synthesized inputs. Digital Threats: Research and Practice (DTRAP), 2(1):4:1– 4:23, March 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3431286.

Glazer:2021:MSL

[GSS21] Amanda K. Glazer, Jacob V.

Spertus, and Philip B. Stark. More style, less work: Cardstyle data decrease risk-limiting audit sample sizes. Digital Threats: Research and Practice (DTRAP), 2(4):32:1–32:15, December 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3457907.

Gajrani:2020:VVV

[GTL^{+}20] Jyoti Gajrani, Meenakshi Tripathi, Vijay Laxmi, Gaurav Somani, Akka Zemmari, and Manoj Singh Gaur. Vulvet: Vetting of vulnerabilities in Android apps to thwart exploitation. Digital Threats: Research and Practice (DTRAP), 1(2): 10:1–10:25, July 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3376121.

Happa:2021:ADS

[HAH^{+}21] Jassim Happa, Ioannis Agrafiotis, Martin Helmhout, Thomas Bashford-Rogers, Michael Goldsmith, and Sadie Creese. Assessing a decision support tool for SOC analysts. Digital Threats: Research and Practice (DTRAP), 2(3):22:1–22:35, September 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3430753.

Happa:2021:DND

[HBRV^{+}21] Jassim Happa, Thomas Bashford-Rogers, Alastair Janse Van Rensburg, Michael Goldsmith, and Sadie Creese. Deception in network defences using unpredictability. Digital Threats: Research and Practice (DTRAP), 2(4):29:1–29:26, December 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3450973.

Huang:2021:IRW

[HZB21] Teng-Chieh Huang, Razieh Nokhbeh Zaeem, and K. Suzanne Barber. Identifying real-world credible experts in the financial domain. Digital Threats: Research and Practice (DTRAP), 2(2): 17:1–17:14, June 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3446783.

Iliou:2021:DAW

[IKT^{+}21] Christos Iliou, Theodoros Kostoulas, Theodora Tsikrika, Vasilis Katos, Stefanos Vrochidis, and Ioannis Kompatsiaris. Detection of advanced Web bots by combining Web logs with mouse behavioural biometrics. Digital Threats: Research and Practice (DTRAP), 2(3):24:1–24:26, September 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3447815.

Jacobs:2021:EPS

[JRE^{+}21] Jay Jacobs, Sasha Romanosky, Benjamin Edwards, Idris Adjerid, and Michael Roytman. Exploit Prediction Scoring System (EPSS). Digital Threats: Research and Practice (DTRAP), 2(3):20:1–20:17, July 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3436242.

Kohlrausch:2020:ASS

[KB20] Jan Kohlrausch and Eugene A. Brin. ARIMA supplemented security metrics for quality assurance and situational awareness. Digital Threats: Research and Practice (DTRAP), 1(1): 6:1–6:21, March 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3376926.

Karim:2020:ADA

[KCH^{+}20] Imtiaz Karim, Fabrizio Cicala, Syed Rafiul Hussain, Omar Chowdhury, and Elisa Bertino. ATFuzzer: Dynamic analysis framework of AT interface for Android smartphones. Digital Threats: Research and Practice (DTRAP), 1(4):23:1– 23:29, December 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3416125.

Kumar:2021:TAG

[KIP21] Rajesh Kumar, Can Isik, and Vir V. Phoha. Treadmill Assisted Gait Spoofing (TAGS): an emerging threat to wearable sensor-based gait authentication. Digital Threats: Research and Practice (DTRAP), 2(3): 23:1–23:17, September 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3442151.

Koot:2020:FNC

[Koo20] Matthijs Koot. Field note on CVE-2019-11510: Pulse connect secure SSL-VPN in The Netherlands. Digital Threats: Research and Practice (DTRAP), 1(2):13:1– 13:7, July 2020. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3382765.

Konstantinovskiy:2021:TAF

[KPBZ21] Lev Konstantinovskiy, Oliver Price, Mevan Babakar, and Arkaitz Zubiaga. Toward automated factchecking: Developing an annotation schema and benchmark for consistent automated claim detection. Digital Threats: Research and Practice (DTRAP), 2(2):14:1– 14:16, June 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3412869.

Laurenza:2020:MTE

[LLM20] Giuseppe Laurenza, Riccardo Lazzeretti, and Luca Mazzotti. Malware triage for early identification of advanced persistent threat activities. Digital Threats: Research and Practice (DTRAP), 1(3):16:1– 16:17, September 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3386581.

Lakhotia:2020:DTR

[LM20] Arun Lakhotia and Leigh Metcalf. Digital threats: Research and practice inaugural issue editorial. Digital Threats: Research and Practice (DTRAP), 1(1):1:1–1:3, March 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3380320.

Lamp:2021:ECA

[LRMZA21] Josephine Lamp, Carlos E. Rubio-Medrano, Ziming Zhao, and Gail-Joon Ahn. Ex-Sol: Collaboratively assessing cybersecurity risks for protecting energy delivery systems. Digital Threats: Research and Practice (DTRAP), 2(3): 21:1–21:23, July 2021. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3428156.

Mainardi:2022:PAC

[MBP22] Nicholas Mainardi, Alessandro Barenghi, and Gerardo Pelosi. Privacy-aware character pattern matching over outsourced encrypted data. Digital Threats: Research and Practice (DTRAP), 3(1):7:1–7:38, March 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3462333.

Metcalf:2021:ESI

[Met21] Leigh Metcalf. Editorial on the special issue on election security. Digital Threats: Research and Practice (DTRAP), 2(4):31:1, December 2021. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3471534.

Mccorry:2021:SVB

[MMT^{+}21] Patrick Mccorry, Maryam Mehrnezhad, Ehsan Toreini, Siamak F. Shahandashti, and Feng Hao. On secure evoting over blockchain. Digital Threats: Research and Practice (DTRAP), 2(4):33:1–33:13, December 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3461461.

Metcalf:2021:EDB

[MS21] Leigh Metcalf and Jonathan M.

Spring. The ecosystem of detection and blocklisting of domain generation. Digital Threats: Research and Practice (DTRAP), 2(3):19:1–19:22, September 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3423951.

Nakatsuka:2021:PPD

[NPT21] Yoshimichi Nakatsuka, Andrew Paverd, and Gene Tsudik. PDoT: Private DNS-over-TLS with TEE support. Digital Threats: Research and Practice (DTRAP), 2(1):3:1–3:22, March 2021. CODEN ????

ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3431171.

Oesch:2022:UPS

[OASD^{+}22] Sean Oesch, Ruba Abu-Salma, Oumar Diallo, Juliane Krämer, James Simmons, Justin Wu, and Scott Ruoti. User perceptions of security and privacy for group chat. Digital Threats: Research and Practice (DTRAP), 3(2):15:1–15:29, June 2022. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3491265.

Osman:2020:SAA

[OMHY20] Tousif Osman, Mohammad Mannan, Urs Hengartner, and Amr Youssef. Securing applications against side-channel attacks through resource access veto. Digital Threats: Research and Practice (DTRAP), 1(4): 22:1–22:29, December 2020. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3416124.

Perdisci:2020:IAS

[PLD^{+}20] Roberto Perdisci, Martina Lindorfer, Adam Doupé, Andrea Lanzi, Alexandros Kapravelos, and Gianluca Stringhini. Introduction to the ACSAC’19 special issue — Part 1. Digital Threats: Research and Practice (DTRAP), 1(4):19e:1– 19e:3, December 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3437251.

Perdisci:2021:IAS

[PLD^{+}21] Roberto Perdisci, Martina Lindorfer, Adam Doupé, Andrea Lanzi, Alexandros Kapravelos, and Gianluca Stringhini. Introduction to the ACSAC’19 special issue — vol. 2. Digital Threats: Research and Practice (DTRAP), 2(1):1:1–1:2, March 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3437253.

Perdisci:2022:IAS

[PLS22] Roberto Perdisci, Martina Lindorfer, and Gianluca Stringhini. Introduction to the AC-SAC’20 special issue. Digital Threats: Research and Practice (DTRAP), 3(2):12:1–12:2, June 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3534708.

Palit:2020:MDO

[PMP20] Tapti Palit, Fabian Monrose, and Michalis Polychronakis. Mitigating data-only attacks by protecting memoryresident sensitive data. Digital Threats: Research and Practice (DTRAP), 1(4):20:1–20:26, December 2020. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3419475.

Rosso:2022:SMT

[RCGA22] Martin Rosso, Michele Campobasso, Ganduulga Gankhuyag, and Luca Allodi. SAIBER-SOC: a methodology and tool for experimenting with security operation centers. Digital Threats: Research and Practice (DTRAP), 3(2):14:1– 14:29, June 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3491266.

Rodriguez:2022:ISI

[RUPT22] Ricardo J. Rodríguez, Xabier Ugarte-Pedrero, and Juan Tapiador. Introduction to the special issue on challenges and trends in malware analysis. Digital Threats: Research and Practice (DTRAP), 3(2):8:1– 8:2, June 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3536319.

Ring:2021:MHB

[RVD^{+}21] John H. Ring, Colin M. Van Oort, Samson Durst, Vanessa White, Joseph P. Near, and Christian Skalka. Methods for host-based intrusion detection with deep learning. Digital Threats: Research and Practice (DTRAP), 2(4):26:1– 26:29, December 2021. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3461462.

Sacher:2020:FFP

[Sac20] Desiree Sacher. Fingerpointing false positives: How to better integrate continuous improvement into security monitoring. Digital Threats: Research and Practice (DTRAP), 1(1):7:1–7:7, March 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3370084.

Sikder:2021:ACA

[SBU21] Amit Kumar Sikder, Leonardo Babun, and A. Selcuk Uluagac. Aegis+: a context-aware platform-independent security framework for smart home systems. Digital Threats: Research and Practice (DTRAP), 2(1): 6:1–6:33, March 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3428026.

Sticha:2022:SOD

[SDA^{+}22] Paul J. Sticha, Tirso E. Diaz, Elise T. Axelrad, Sean D. Vermillion, and Dennis M. Buede. Simulating organizational data from redacted input for inference enterprise modeling. Digital Threats: Research and Practice (DTRAP), 3(1):4:1–4:30, March 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3457910.

Shukla:2020:TUI

[SKM^{+}20] Diksha Shukla, Partha Pratim KUNDU, Ravichandra Malapati, Sujit Poudel, Zhanpeng Jin, and Vir V. Phoha. Thinking unveiled: an inference and correlation model to attack EEG biometrics. Digital Threats: Research and Practice (DTRAP), 1(2):9:1–9:29, July 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3374137.

Samtani:2021:ICT

[SLBC21] Sagar Samtani, Weifeng Li, Victor Benjamin, and Hsinchun Chen. Informing cyber threat intelligence through dark Web situational awareness: The AZSecure hacker assets portal. Digital Threats: Research and Practice (DTRAP), 2(4):27:1– 27:10, December 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3450972.

Scofield:2020:AML

[SMK20] Daniel Scofield, Craig Miles, and Stephen Kuhn. Automated model learning for accurate detection of malicious digital documents. Digital Threats: Research and Practice (DTRAP), 1(3):15:1–15:21, September 2020. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3379505.

Sanchez-Rola:2020:CPT

[SRBS20] Iskander Sanchez-Rola, Davide Balzarotti, and Igor Santos. Cookies from the past: Timing server-side request processing code for history sniffing. Digital Threats: Research and Practice (DTRAP), 1(4):24:1– 24:24, December 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3419473.

Steinhauser:2020:DTI

[ST20] Antonín Steinhauser and Petr Tuma. Database traffic interception for graybox detection of stored and context-sensitive XSS. Digital Threats: Research and Practice (DTRAP), 1(3): 17:1–17:23, September 2020. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3399668.

Sun:2022:CRV

[Sun22] Sheng Sun. A chosen random value attack on WPA3 SAE authentication protocol. Digital Threats: Research and Practice (DTRAP), 3(2):16:1–16:8, June 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3468526.

Tien:2020:MLF

[TCBK20] Chin-Wei Tien, Shang-Wen Chen, Tao Ban, and Sy-Yen Kuo. Machine learning framework to analyze IoT malware using ELF and opcode features. Digital Threats: Research and Practice (DTRAP), 1(1):5:1–5:19, March 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3378448.

Ukrop:2020:WYT

[UKM20] Martin Ukrop, Lydia Kraus, and Vashek Matyas. Will you trust this TLS certificate?: Perceptions of people working in IT (extended version). Digital Threats: Research and Practice (DTRAP), 1(4):25:1– 25:29, December 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3419472.

Usui:2021:ARE

[UOI^{+}21] Toshinori Usui, Yuto Otsuki, Tomonori Ikuse, Yuhei Kawakoya, Makoto Iwamura, Jun Miyoshi, and Kanta Matsuura. Automatic reverse engineering of script engine binaries for building script API tracers. Digital Threats: Research and Practice (DTRAP), 2(1): 5:1–5:31, March 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3416126.

VanDerHam:2021:TBU

[Van21] Jeroen Van Der Ham. Toward a better understanding of “cybersecurity”. Digital Threats: Research and Practice (DTRAP), 2(3):18:1–18:3, September 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3442445.

VanderToorn:2020:LBH

[VS20] Olivier Van der Toorn and Anna Sperotto. Looking beyond the horizon: Thoughts on proactive detection of threats. Digital Threats: Research and Practice (DTRAP), 1(1):4:1– 4:13, March 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3373639.

Valecha:2021:FNS

[VSV^{+}21] Rohit Valecha, Srikrishna Krishnarao Srinivasan, Tejaswi Volety, K. Hazel Kwon, Manish Agrawal, and H. Raghav Rao. Fake news sharing: an investigation of threat and coping cues in the context of the Zika virus. Digital Threats: Research and Practice (DTRAP), 2(2):16:1–16:16, June 2021. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3410025.

Williams:2022:RIA

[WASC22] Adam D. Williams, Shannon N.

Abbott, Nathan Shoman, and William S. Charlton. Results from invoking artificial neural networks to measure insider threat detection & mitigation. Digital Threats: Research and Practice (DTRAP), 3(1):3:1–3:20, March 2022. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3457909.

Waked:2020:SST

[WMY20] Louis Waked, Mohammad Mannan, and Amr Youssef. The sorry state of TLS security in enterprise interception appliances. Digital Threats: Research and Practice (DTRAP), 1(2):8:1–8:26, July 2020. CO-DEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3372802.

Zimmer:2022:IDN

[ZBF22] Ephraim Zimmer, Christian Burkert, and Hannes Federrath. Insiders dissected: New foundations and a systematisation of the research on insiders. Digital Threats: Research and Practice (DTRAP), 3(1): 2:1–2:35, March 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3473674.

Zhou:2020:FNE

[ZJPZ20] Xinyi Zhou, Atishay Jain, Vir V. Phoha, and Reza Zafarani. Fake news early detection: a theory-driven model. Digital Threats: Research and Practice (DTRAP), 1(2):12:1– 12:25, July 2020. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3377478.

Zafarani:2021:IRT

[ZLPA21] Reza Zafarani, Huan Liu, Vir V.

Phoha, and Javad Azimi. Inroduction on recent trends and perspectives in fake news research. Digital Threats: Research and Practice (DTRAP), 2(2):13:1–13:3, March 2021. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3448634.

Zheng:2022:UDM

[ZYW22] Panpan Zheng, Shuhan Yuan, and Xintao Wu. Using Dirichlet marked Hawkes processes for insider threat detection. Digital Threats: Research and Practice (DTRAP), 3(1):5:1– 5:19, March 2022. CODEN ???? ISSN 2692-1626 (print), 2576-5337 (electronic). URL

https://dl.acm.org/doi/10.
1145/3457908.